# Fantastic Targets for Concept Erasure in Diffusion Models and Where To Find Them

### v2

@@ 1 Introduction
The widespread accessibility of text-to-image generation models has introduced significant risks such as the generation of harmful content, copyright infringements, and biases due to the exposure of models to undesirable concepts during training. Initial attempts to address these concerns focus on dataset filtering during the training phase (StabilityAI, 2022), post-generating filtering (Rando et al., 2022) or inference guiding (Schramowski et al., 2023b). While dataset filtering often requires costly retraining, post-processing and inference-based methods can be easily bypassed (Yang et al., 2024). Recently, concept erasure (Bui et al., 2024b; a; Gandikota et al., 2023; Orgad et al., 2023; Zhang et al., 2023; Kumari et al., 2023), which aims to directly remove the concept from the model’s parameters without the need for complete retraining, has emerged as a promising alternative.
Concept erasure methods can be broadly categorized into two approaches: output-based and attention-based. Output-based methods aim to neutralize the output associated with undesirable concepts (Bui et al., 2024b; a; Gandikota et al., 2023; Wu et al., 2024), while attention-based methods modify the attention scores of these concepts in the cross-attention layers of the model (Zhang et al., 2023; Orgad et al., 2023; Gandikota et al., 2024; Lu et al., 2024; Lyu et al., 2024). Despite their differences, both approaches share a common principle: mapping undesirable concepts to a fixed, generic target, such as “a photo” or an empty text prompt. Although these methods have shown success in erasing undesirable concepts, they do not consider how the choice of target concepts affects both the effectiveness of erasure and the preservation of benign concepts.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our method in erasing various concepts from the foundation model, including object-related concepts, artistic styles, and NSFW attributes. We compare our approach with the state-of-the-art erasure methods including AP (Bui et al., 2024b), ESD (Gandikota et al., 2023), UCE (Gandikota et al., 2024), CA (Kumari et al., 2023), and MACE (Lu et al., 2024). Our experiments follow the same setup as in Bui et al. (2024b); Gandikota et al. (2023; 2024). Specifically, we use Stable Diffusion (SD) version 1.4 as the foundation model, and fine-tune the model for 1000 steps with a batch size of 1, using the Adam optimizer with a learning rate of $\alpha=10^{-5}$.
@@ Acknowledgements
The Commonwealth of Australia (represented by the Advanced Strategic Capabilities Accelerator) supports this research through a Defence Science Partnerships agreement. Dinh Phung further acknowledges the support from the Australian Research Council (ARC) Discovery Project DP250100262. The authors would like to thank the anonymous reviewers for their insightful feedback and valuable suggestions, which have significantly enhanced the quality of this work.
@@ Appendix A Related Work
Adversarial Concepts.
The definition of adversarial concepts- those most affected by changes in model parameters when a concept is erased- was first introduced in (Bui et al., 2024b). Building on empirical observations that erasing different concepts impacts the remaining ones in various ways and that preserving a neutral concept alone is insufficient to maintain the model’s capabilities, the authors propose a method for preserving sensitive concepts while erasing undesirable ones. This approach is similar to the adversarial training principle (Goodfellow et al., 2014; Bui et al., 2022; 2021; 2020) which has been successfully applied to improve the robustness of models against adversarial attacks. Our work is inspired by the idea of adversarial concepts but with several key differences: ❶ Instead of focusing on the to-be-preserved concept, we investigate the impact of the target concept on the remaining concepts, offering new insights into the selection and importance of the target concept in the erasure problem. ❷ We propose an improved optimization method that more accurately identifies adversarial concepts, explicitly excluding those identical to the to-be-erased concept—an aspect not considered in (Bui et al., 2024b). ❸ For the first time, we introduce the concept of a concept graph and its properties, providing a deeper understanding of concept relationships in the latent space.
@@ D.4 Evaluation on Scalability
Table 7 shows the breakdown of the concepts (to-be-erased and to-be-preserved-similar) used in the experiment.

### v3

@@ 1 Introduction
The widespread accessibility of text-to-image generation models has introduced significant risks such as the generation of harmful content, copyright infringements, and biases due to the exposure of models to undesirable concepts during training. Initial attempts to address these concerns focus on dataset filtering during the training phase (StabilityAI, 2022), post-generating filtering (Rando et al., 2022) or inference guiding (Schramowski et al., 2023b). While dataset filtering often requires costly retraining, post-processing and inference-based methods can be easily bypassed (Yang et al., 2024). Recently, concept erasure (Bui et al., 2024a; b; Gandikota et al., 2023; Orgad et al., 2023; Zhang et al., 2023; Kumari et al., 2023), which aims to directly remove the concept from the model’s parameters without the need for complete retraining, has emerged as a promising alternative.
Concept erasure methods can be broadly categorized into two approaches: output-based and attention-based. Output-based methods aim to neutralize the output associated with undesirable concepts (Bui et al., 2024a; b; Gandikota et al., 2023; Wu et al., 2024), while attention-based methods modify the attention scores of these concepts in the cross-attention layers of the model (Zhang et al., 2023; Orgad et al., 2023; Gandikota et al., 2024; Lu et al., 2024; Lyu et al., 2024). Despite their differences, both approaches share a common principle: mapping undesirable concepts to a fixed, generic target, such as “a photo” or an empty text prompt. Although these methods have shown success in erasing undesirable concepts, they do not consider how the choice of target concepts affects both the effectiveness of erasure and the preservation of benign concepts.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our method in erasing various concepts from the foundation model, including object-related concepts, artistic styles, and NSFW attributes. We compare our approach with the state-of-the-art erasure methods including AP (Bui et al., 2024a), ESD (Gandikota et al., 2023), UCE (Gandikota et al., 2024), CA (Kumari et al., 2023), and MACE (Lu et al., 2024). Our experiments follow the same setup as in Bui et al. (2024a); Gandikota et al. (2023; 2024). Specifically, we use Stable Diffusion (SD) version 1.4 as the foundation model, and fine-tune the model for 1000 steps with a batch size of 1, using the Adam optimizer with a learning rate of $\alpha=10^{-5}$.
@@ Acknowledgements
This work was supported by the Australian Defence Science and Technology (DST) Group through the Next Generation Technology Fund (NGTF) scheme and the Department of Defence, Australia, via the Advanced Strategic Capabilities Accelerator (ASCA) program. Dinh Phung further acknowledged the support from the Australian Research Council (ARC) Discovery Project DP250100262. The authors would like to thank the anonymous reviewers for their insightful feedback and valuable suggestions, which have significantly enhanced the quality of this work.
@@ D.4 Evaluation on Scalability
Table LABEL:tab:concept_space shows the breakdown of the concepts (to-be-erased and to-be-preserved-similar) used in the experiment.